OpenStudy (anonymous):

Is (R/{0}, *) isomorphic to (R, +)?

OpenStudy (anonymous):

Of course not :)

OpenStudy (anonymous):

But you're probably not gonna take my word for it, and you require an explanation, am I right? ^.^

OpenStudy (anonymous):

>.> Anyway... Let's let G = <R\{0} , *> H=<R,+> And suppose \[\Large G\cong H\]

OpenStudy (anonymous):

Well, since they're isomorphic, there's a function \[\Large \phi:G\rightarrow H\]such that for any two elements g_1 and g_2 in G \[\Large \phi(g_1 \times g_2)=\phi(g_1)+\phi(g_2)\]

OpenStudy (anonymous):

Also, since it's an isomorphism, identity in G must be mapped to the identity in H. \[\Large \phi(1)=0\]

OpenStudy (anonymous):

So, let's get started :) We don't know what \[\Large \phi(-1) \]is, but let's make it a real number x. \[\Large \phi(-1)=x\] Now, because of the isomorphism \[\Large \phi(-1 \times -1)=\phi(-1)+\phi(-1)=x + x=2x\]

OpenStudy (anonymous):

But \[\Large -1 \times -1=1\]implying\[\Large 2x=\phi(-1\times-1)=\phi(1)=0\]Therefore \[\Large 2x=0\]\[\Large x=0\]

OpenStudy (anonymous):

Then that means \[\Large \phi(-1)=x=0\] Which means \[\Large \phi(-1)=0=\phi(1)\] Means \[\LARGE \phi:G\rightarrow H\]cannot be bijective, and thus, cannot be an isomorphism Case closed ^.^